OpenStudy (anonymous):

Solve 2(3x – 4) = 3x + 7. You must show all work.

OpenStudy (anonymous):

2(3x-4) = 3x + 7 6x-8 = 3x + 7 add 8 to both ides and subtract 3x from both sides, you get: 6x-3x = 8+7 3x = 15 divide both side by 3: x = 5

OpenStudy (anonymous):